OpenStudy (anonymous):

Solve: −2(x − 1) = −16

OpenStudy (anonymous):

okay so distribute -2(x-1)

OpenStudy (anonymous):

-2x-2?

OpenStudy (anonymous):

not quite

OpenStudy (bahrom7893):

^correct :D

OpenStudy (bahrom7893):

lol robz the fact that your response was before mine saved me from embarrassment

OpenStudy (anonymous):

-2x+2

OpenStudy (bahrom7893):

^Now that is correct

OpenStudy (anonymous):

@bahrom7893 ;) @shahbaaz786 correct

OpenStudy (bahrom7893):

double negative = positive

OpenStudy (anonymous):

okay now set that equal to -16

OpenStudy (anonymous):

7?

OpenStudy (anonymous):

not quite -2x + 2 = -16 what is -2x +2 - 2 = -16 - 2 ?

OpenStudy (anonymous):

-2x=-18

OpenStudy (anonymous):

correct, now solve for x

OpenStudy (anonymous):

9?

OpenStudy (anonymous):

there ya go